### v1

# Enhancing LLM-based Test Generation for Hard-to-Cover Branches via Program Analysis

Abstract
Automatic test generation plays a critical role in software quality assurance. While the recent advances in Search-Based Software Testing (SBST) and Large Language Models (LLMs) have shown promise in generating useful tests, these techniques still struggle to cover certain branches. Reaching these hard-to-cover branches usually requires constructing complex objects and resolving intricate inter-procedural dependencies in branch conditions, which poses significant challenges for existing techniques. In this work, we propose TELPA, a novel technique aimed at addressing these challenges. Its key insight lies in extracting real usage scenarios of the target method under test to learn how to construct complex objects and extracting methods entailing inter-procedural dependencies with hard-to-cover branches to learn the semantics of branch constraints. To enhance efficiency and effectiveness, TELPA identifies a set of ineffective tests as counter-examples for LLMs and employs a feedback-based process to iteratively refine these counter-examples. Then, TELPA integrates program analysis results and counter-examples into the prompt, guiding LLMs to gain deeper understandings of the semantics of the target method and generate diverse tests that can reach the hard-to-cover branches. Our experimental results on 27 open-source Python projects demonstrate that TELPA significantly outperforms the state-of-the-art SBST and LLM-based techniques, achieving an average improvement of 31.39% and 22.22% in terms of branch coverage.
I Introduction
Automatic test generation holds significant importance in software quality assurance, which aims to efficiently detect software bugs by automatically covering various behaviors of the software under test. Recently, many test generation techniques have been proposed, among which Search-Based Software Testing (SBST) [1, 2, 3] is one of the most widely-studied categories due to its effectiveness. Typically, SBST techniques leverage heuristic search algorithms to explore the vast test space by mutating already-generated tests with the guidance of maximizing code coverage. While receiving extensive attention, they are still unable to reach specific parts of code. For instance, they often fail to generate tests to cover the branches that can be triggered only by some scenario-specific values (e.g., specific file names, version strings) due to the lack of the ability of understanding code semantics [4].
The recent advancement of Large Language Models (LLMs) provides new opportunities for tackling these problems, and several LLM-based techniques have been proposed to generate effective test cases [4, 5]. These techniques utilize the code comprehension ability of LLMs by incorporating the source code of the target method to be tested and some contextual information (such as other methods around the target method) into the prompt. While the limitation of SBST can be relieved to some degree with the help of LLMs, the automatically generated tests still tend to be generic and only able to cover the branches without complicated constraints [6]. That is, many hard-to-cover branches are still left for manual test design based on experts’ domain knowledge, leading to significant costs. An automated approach which can further enhance the coverage of these hard-to-cover branches can largely reduce the testing effort.
In practice, two major challenges stand out when creating such approaches. First, some branch constraints involve objects with complex construction processes. Specifically, the construction of some objects relies on other (complex) types of objects, ultimately necessitating a specific sequence of constructor invocations. As indicated by a previous study [5], the tests generated by LLMs exhibit a relatively low compilation success rate (about 39%), highlighting a deficiency in the ability of LLMs in this regard. Hence, state-of-the-art LLM-based techniques (e.g., CODAMOSA [4]) fail to produce valid objects for such scenarios, let alone the objects with specific attribute values required by branch constraints. Second, some branch constraints entail intricate inter-procedural dependencies. That is, the outcome of certain conditions within the constraints are determined by the execution results of a series of method invocations. In such scenarios, merely providing source code of the target method or some coarse-grained contextual information (the common practice of existing LLM-based techniques) is insufficient to enable LLMs to comprehend the semantics conveyed by a sequence of invoked methods, thereby struggling to generate effective tests.
To tackle these challenges, we propose a novel LLM-based test generation technique (called TELPA – TEst generation via Llm and Program Analysis), which focuses on reaching hard-to-cover branches via program-analysis-enhanced prompting. Given that complex objects in branch constraints are often passed to the target method as parameters, TELPA addresses the first challenge by gathering invocations of the target method within the software module. These existing invocations might contain the whole process of constructing these objects. Specifically, TELPA conducts backward method-invocation analysis, starting from each invocation of the target method, to extract all sequences of method invocations with the target method as the endpoint. Then, TELPA prompts LLMs to generate tests by using different paths to enter the target method. By exploring different entry paths, TELPA not only facilitates the construction of valid objects but also captures various usage scenarios in practice (that may produce specific attribute values required by branch constraints), thus enhancing test diversity and improving coverage.
To address the second challenge arising from complicated inter-procedural dependencies in branch constraints, TELPA performs forward method-invocation analysis. This analysis begins from each variable and invoked method in the constraints for a target branch, aiming to extract all associated methods. By incorporating the source code of these methods in their invocation order as the prompt, TELPA can elicit LLMs to gain a deeper understanding of the semantics pertaining to the target branch, thereby generating effective tests. This provides precise contextual information concerning the target branch and it is more effective than directly supplying the source code of all methods within the software module to LLMs, as the latter could introduce excessive noise irrelevant to the target branch and confuse LLMs during test generation.
Due to the non-negligible cost of utilizing LLMs, we define the usage scenario of TELPA as its activation when the existing (lightweight) test generation tool (e.g., Pynguin [1]) fails to increase test coverage within a certain timeframe. In other words, TELPA is only employed for hard-to-cover branches to ensure cost-effectiveness. To enhance efficiency further, TELPA samples a diverse set of already-generated tests as counter-examples. By incorporating these counter-examples into the prompt, TELPA can instruct LLMs to generate tests that diverge from them, as these counter-examples have been recognized as ineffective for these hard-to-cover branches. In particular, the overall test generation process with TELPA is structured as a feedback-based process. This design allows for updating counter-examples based on the most recent coverage results, thereby enhancing test generation effectiveness.
We conducted an extensive evaluation on 27 open-source Python projects, which have been widely used in previous studies [4, 1]. Our results show that TELPA significantly outperforms both Pynguin (a state-of-the-art SBST technique) and CODAMOSA (a state-of-the-art LLM-based technique). On average, TELPA achieves 31.39% and 22.22% higher branch coverage than them respectively, given the same testing time budget. Our ablation study also confirms the contribution of each main component in TELPA, including backward method-invocation analysis for relieving the challenge arising from complex object construction, forward method-invocation analysis for relieving the challenge arising from intricate inter-procedural dependencies, counter-examples sampling and coverage-based feedback for improving the overall testing efficiency. Our investigation into various configurations of TELPA confirms that our designed usage scenario for TELPA strikes a balance between effectiveness and efficiency and TELPA consistently achieves higher coverage than other settings.
The main contributions of our study are as follows.
We introduce TELPA, a novel LLM-based test generation technique specifically designed to enhance the coverage of hard-to-cover branches through program-analysis-enhanced prompting.
TELPA can well address the challenges arising from complex object construction and intricate inter-procedural dependencies by incorporating both backward and forward method-invocation analysis within TELPA.
The feedback-based test generation process in TELPA improves the efficiency and effectiveness of test generation by guiding LLMs with diverse counter-example tests.
We present an extensive evaluation of TELPA across 27 open-source projects, demonstrating its effectiveness in comparison to a state-of-the-art SBST technique and a state-of-the-art LLM-based one.
II Motivating Example
In this section, we demonstrate the limitations of the state-of-the-art techniques Pynguin [1] and CODAMOSA [4] with two real examples to motivate our study.
Figure 1 shows the simplified code snippet of the target method set_definition from the typesystem project111https://github.com/encode/typesystem. With a time budget of 20 minutes, both Pynguin and CODAMOSA fail to reach line 6, namely the condition for the if statement in line 5 never returns true. In fact, to make the condition true, the generated test needs to first construct an Array object (line 3), and then assign a proper value to its items attribute. More specifically, items should be either a tuple or a list of Field objects (line 4), according to the type hints from the constructor of Array (line 13). Constructing such a sophisticated Array object is certainly challenging for automatic test generation techniques, Pynguin and CODAMOSA are no exception.
Another type of hard-to-cover branches are caused by complicated inter-procedural dependencies, as illustrated in Figure 2. In this case, our target method is is_public_family from the apimd project222https://github.com/KmolYuan/apimd. Both Pynguin and CODAMOSA are unable to guide the condition in line 3 to return true. The challenge here is that the outcome of the if condition depends on the method is_magic, which is defined elsewhere in the project and might further invoke other methods. Those complicated inter-procedural dependencies are often hard to understand and analyze for test generation techniques.
There are several reasons which might contribute to the inability of the state-of-the-art techniques to handle branches with complex objects or inter-procedural dependencies.
First, during the test generation process, existing techniques [1, 4] normally try to construct minimal or over-simplified objects, without taking into account the real usage scenarios involving specific attributes. Second, for the inter-procedural dependencies, state-of-the-art techniques [1, 4] are not able to recursively analyze how dependent methods interact with variables from the target method, thus failing to interpret the concrete branch constraints. Third, existing LLM-based techniques [4, 5] typically provide the whole module as the context for LLMs to generate tests. In reality, many parts of code is irrelevant to the target method, which can only introduce noise to LLMs. Moreover, a prompt can easily exceed the input length limit if the information of the whole module is incorporated, resulting in undesired prompt truncation and critical information loss.
Based on these potential reasons behind the unsatisfactory performance of existing techniques, we propose TELPA, an LLM-based test generation technique leveraging program-analysis-enhanced prompting. TELPA will 1) conduct backward method-invocation analysis to learn real usage scenarios of objects and overcome the difficulty of constructing complex objects, 2) conduct forward method-invocation analysis to understand the behavior of complicated inter-procedural dependencies, and 3) only supply relevant information to LLMs for test generation, avoiding information interference.
III Approach
The overall workflow of TELPA is illustrated in Figure 3. First, TELPA leverages existing test generation tools to generate a set of tests, which can reach those easy-to-cover branches. For ease of presentation, we call this step preceding test generation. Then, for the methods containing the branches that are still not covered, TELPA performs two types of program analyses, including backward method-invocation analysis to retrieve the method invocation sequences for complex object construction and forward method-invocation analysis to retrieve all associated methods for inter-procedural dependencies. To improve the testing efficiency, TELPA samples existing generated tests as counter-examples to guide LLMs to generate different tests, as they are shown ineffective to cover specific branches. These counter-examples, together with the program analysis results, are incorporated into the prompt for LLMs to generate new tests. The new tests will be executed and added into existing tests for future iterations.
III-A Preceding Test Generation with Existing Tools
As our goal is to reach those hard-to-cover branches, TELPA outsources the initial testing task (for those easy-to-cover branches) to the existing test generation tools. This helps to improve the cost-effectiveness of the overall testing process due to the cost of utilizing LLMs. Specifically, for a project under test, TELPA runs selected state-of-the-art test generation tools until they are unable to cover new branches within a pre-defined timeframe. When executing these generated tests, their corresponding coverage information is also recorded. It is worth noting that TELPA is in theory language-independent and not restricted to specific tools for generating initial tests. In this section, we mainly describe the idea behind TELPA, and the implementation details will be presented in Section IV-A1.
III-B Method-Invocation Analyses
TELPA first collects the methods containing branches not covered by the preceding testing process, and then performs backward and forward method-invocation analyses on these methods to gather information needed for handling complex object construction and inter-procedural dependencies. This process is done in ascending order of coverage achieved for the methods, as we assume that methods with the lower coverage are more difficult to tackle and there is larger room for coverage improvement. Unlike existing LLM-based techniques (e.g., CODAMOSA [4]) that try to incorporate as much source code from the module as possible into prompts, we will only feed LLMs with the methods retrieved from method-invocation analyses in order to reduce noise introduced by irrelevant code and alleviate computational burden of LLMs.
III-B1 Backward Analysis
The goal of the backward analysis is to extract method-invocation sequences ending with the target method. These sequences represent diverse real usage scenarios of the target method in the project and have a high chance to capture the whole process of constructing objects used in the target method. For example, a method might create an object o and invoke the target method v with o as an argument. A backward method-invocation analysis would allow us to trace back to the object construction process.
To achieve this goal, TELPA first constructs a method call graph within the module hosting the target method, and then extracts all paths terminating at the target method as the method invocation sequences.
Call Graph Construction. A directed call graph is constructed to represent the method invocation relationships among methods in the module hosting the target method. The nodes in the graph represent the methods in the module extracted by traversing the abstract syntax tree (AST), and the edges represent invocation relationships between pairs of methods. Figure 4 illustrates an example of a constructed call graph. Given the simplified code snippet, the call graph will contain nodes method1 (denoted by ) from RelevantClass1, method2 (denoted by ) and method3 (denoted by ) from RelevantClass2. Clearly, is called by and , while is only called by .
For each extracted method, TELPA also records the associated classes (the parent class and classes utilized within the method), as well as their declarations and constructors. This information is important because extracted methods might belong to or interact with variables/methods from multiple different classes. Understanding where these elements come from is critical for LLMs to construct valid objects.
Sequence Extraction. With the directed call graph, TELPA proceeds to extract sequences of method invocations for the target method. Specifically, TELPA identifies all the entry nodes from the graph and employs the Depth-First Search (DFS) strategy to gather all the paths originating from entry nodes and terminating at the target method. If a cycle is encountered during path extraction, the traversal of that path will stop to prevent creating infinite paths. It is worth noting that for all public methods inside a class, we consider each method itself as a complete path whose entry method is also the target method. Going back to the example in Figure 4, for the target method , three paths will be extracted: , and . As can be seen here, there might be multiples call paths from a same entry method to the target method (e.g., and ).
To alleviate computational load, TELPA employs a straightforward filtering strategy to condense these sequences with the same entry method to the target method. That is, only the shortest path is kept. This strategy is adopted due to such a shorter and more direct call paths are easier for LLMs to understand and learn from. In our example, the path will be selected (that is, will be filtered out).
III-B2 Forward Analysis
As illustrated in Section II, it is challenging for test generation techniques to reach certain branches when the outcome of the branch condition depends on other invoked methods. To address it, TELPA performs a forward method-invocation analysis to recursively collect all methods that might be relevant to determining which branch to execute. The extracted methods provide rich information for interpreting what really entails in branch conditions. Below, we describe the forward analysis process in detail with the example presented in Figure 5.
To extract the set of all relevant methods for a target method (example in Figure 5), TELPA first identifies branch conditions and captures all variables (var1, param1) and methods (method2) these condition expressions contain. All involved methods (method2) are added to . For each involved variable, TELPA recursively extracts relevant assignment statements. In our example, the value of var1 is determined by method1, which takes an argument var0 determined by method0. TELPA considers both assignments are relevant to var1 and adds the involved methods (method0 and method1) to . In this way, all methods that can impact the outcome of branch conditions are retrieved and a precise search scope is identified to understand the behavior of condition expressions.
III-C Counter-Examples Sampling
With the preceding testing process described in Section III-A, we have acquired a set of tests that fail to reach specific hard-to-cover branches. These existing tests can be used as counter-examples to guide LLMs to generate divergent tests, further enhancing the overall test efficiency. However, incorporating all existing generated tests into the prompt is impractical, which will not only impose significant computational burden on LLMs but also easily exceeds the input length limit of the prompt. To mitigate this, TELPA samples a diverse set of counter-example tests, striking a balance between effectiveness and efficiency. As our goal is to reach hard-to-cover branches, it is essential to let LLM understand which branches have already been covered by existing tests. Meanwhile, as mentioned before, we want to have a small (preferably minimal) set of tests as counter-examples.
To achieve this, TELPA employs a coverage-based approach. For each method invocation sequence of a target method obtained from the backward analysis, TELPA collects all the existing tests that invoke the first method of the sequence (denoted by ) as the candidate tests (denoted by ). We only consider the first method of the sequence as it will in the end propagate to the target method through the chain of method calls. From , TELPA first selects the test that achieves the highest coverage for the target method , then picks the next tests that can achieve the highest incremental coverage. This is an iterative process until no more test can increase the coverage. In this way, the minimal set of tests can be acquired to reach all already-covered branches. As can be seen, the counter-examples are constructed for each method invocation sequence, this is because in the later stage of test generation, each time we will only feed one method invocation sequence to the LLM to alleviate the computational load and avoid exceeding the input length limit.
III-D Feedback-based Test Generation Process with LLMs
TELPA incorporates program analysis results and counter-examples into the prompt for the LLM to generate new tests. Previous studies have shown large performance improvement for LLMs through Chain-of-Thought (CoT) [7, 8]. Therefore, we also adopt a typical CoT strategy and divide the process into two stages. Note that TELPA can also adopt different CoT techniques, as the contribution of TELPA lies in improving the prompting approach instead of designing CoT techniques.
In the first stage, TELPA constructs the context of the target method with two sources: 1) methods in a selected method-invocation sequence from backward analysis, and 2) the associated methods from forward analysis, along with the declarations and constructors of their associated classes. To select the method-invocation sequence for in each iteration, TELPA first uses the sequence (i.e., the entry method is itself), as this is the most straightforward sequence for the LLM to invoke and understand the semantics of . If this sequence does not exist (e.g., the target method is not a public method), or the generated tests with it in the second stage do not cover new branches (note that our testing process is iterative with feedback), TELPA randomly selects an unused sequence. The context information is then fed to the LLM for summarizing the functionality of , enabling the LLM to gain deep understandings of the method semantics. The test generation process for terminates when no sequence is left.
In the second stage, TELPA integrates the contents of the sampled counter-examples corresponding to the method invocation sequence into the prompt, along with a description: “these counter-examples enter the target method via the selected sequence of method invocations.” Such a description can facilitate the understanding of the LLM on the intent of the generation, namely to cover different branches in the target method following the method invocation sequence. The LLM is then instructed to generate divergent tests.
The new test generated by the LLM is executed and added to the existing tests along with its coverage information. The used counter-examples are removed to ensure the diversity in future counter-example selection. This type of feedback helps TELPA generate more tests with different behaviors.
IV Evaluation
To evaluate the performance of TELPA, we formulate the following research questions (RQs).
RQ1: To what extent can TELPA improve code coverage compared to the state-of-the-art techniques? The goal of TELPA is to reach hard-to-cover branches, and this RQ verifies its ability in this regard.
RQ2: How does each of the main components in TELPA affect the performance? This RQ leads to an ablation study to examine how TELPA’s main components (i.e., method-invocation analysis, counter-example sampling, coverage-based feedback) impact the performance.
RQ3: How does the adoption of preceding test generation tools and LLMs impact the performance of TELPA? Our TELPA outsources initial test generation to existing tools, and utilizes an LLM to generate new tests. In this RQ, we examine how the adoption of existing tools and the choice of the LLM can impact the performance.
IV-A Experimental Setup
IV-A1 Implementation of TELPA
While TELPA is not limited to specific programming languages, in this experiment, we target test generation for Python programs, given the popularity of Python and the existing Python benchmarks [4]. TELPA is also developed in Python 3.10 with over 6k lines of code. The LLM Phind-CodeLlama-34B-v2 [9] from Hugging Face [10] is employed for test generation, as it is one of the most effective open-source code LLMs. PyTorch 1.13 [11], Hugging Face Transformers 4.35.2 [12], and fastchat 0.2.30 [13] are used to run the LLM locally. Following the existing work [14, 15], we set the temperature to 1, enabling TELPA to obtain results from the LLM with a high degree of confidence.
To reduce the resource consumption, TELPA outsources initial test generation to state-of-the-art techniques. In our study, we adopted two test generation tools separately: a state-of-the-art SBST technique Pynguin [1] and a state-of-the-art LLM-based technique CODAMOSA [4]. That is, we constructed two instances of TELPA: Pynguin-based TELPA and CODAMOSA-based TELPA. The adoption of these two different types of preceding test generation techniques ensures the conclusion generalizability. More information regarding Pynguin and CODAMOSA will be given in Section IV-A2.
As mentioned before, TELPA is activated when the preceding tool fails to increase coverage within a pre-defined timeframe. As the previous work [4], we conducted a pilot study on a small test benchmark to optimize the performance of TELPA. The timeframe is set to 2min based on the observed results, namely TELPA will take over the test generation task if Pynguin or CODAMOSA cannot increase the coverage for two minutes. This choice is further discussed in Section IV-E.
IV-A2 Baselines for RQ1
To answer RQ1, we compared TELPA with Pynguin and CODAMOSA, the techniques used for initial test generation in TELPA.
Pynguin [1] is the state-of-the-art SBST technique for dynamically typed programming languages like Python. It initiates the testing process by taking the Python code under test as input, and employs one of search-based algorithms (including MIO [16, 17], MOSA [18], and DynaMOSA [18]) to generate new tests by mutating values and statements of a seed test. Here, we used Pynguin 0.19.0 with DynaMOSA as this algorithm has been demonstrated to be the most effective one among all supported algorithms in Pynguin [1].
CODAMOSA [4] is the state-of-the-art LLM-based test generation tool, which was proposed for improving SBST techniques. It prompts LLMs to generate seed tests for Pynguin in order to help SBST escape the local optimum. More specifically, when Pynguin falls into the local optimum, CODAMOSA switches to invoke LLMs to generate new seed tests by incorporating as much source code of the module under test into the prompt as possible, and then switches back to Pynguin for running on these new seed tests, until the given testing budget is reached. CODAMOSA originally employed Codex as the LLM for generating seed tests. However, as the Codex service has been shut down [19], we employed Phind-CodeLlama-34B-v2 as the LLM for CODAMOSA to have a fair comparison with our technique.
One of the main contributions of TELPA is the novel prompting method. Unlike CODAMOSA that incorporates as much source code of the module under test as possible for prompting, TELPA only feed part of the code which is highly relevant to branch constraints to LLMs. The comparison with the LLM-based CODAMOSA cannot clearly demonstrate the impact of this new prompting method due to the different LLM invocation workflows. Therefore, we constructed a new LLM-based baseline (called BasicPrompt), which replaces our novel prompting method with the one used CODAMOSA (i.e., incorporating as much code as possible). In fact, this traditional prompting approach is commonly used to generate tests in practice [4, 5]. Since in our study, both Pynguin and CODAMOSA are used as preceding test generation tools, we created two instances of BasicPrompt for comparisons: Pynguin-based BasicPrompt and CODAMOSA-based BasicPrompt.
To sum up, we have two sets of comparisons: 1) TELPA vs. Pynguin vs. BasicPrompt and 2) TELPA vs. CODAMOSA vs. BasicPrompt. These comparisons aim to investigate whether TELPA can improve the existing test generation tools (Pynguin and CODAMOSA) and examine whether our novel prompting method can outperform the widely-used prompting method regardless of the preceding test generation tools.
In our study, we did not compare TELPA with the other LLM-based test generation techniques such as ChatUnitTest [20] and ChatTester [5]. The reason is that these approaches leveraged similar prompting method as CODAMOSA. That is, they also feed LLMs with as much available information as possible, such as target methods and other methods around target methods. Particularly, their main novelty which contributes the performance improvement is the post-processing strategies adopted to repair invalid tests generated by LLMs. Those test-repair strategies are orthogonal to the contribution of TELPA, which can be also integrated into TELPA to further improve its overall test effectiveness.
IV-A3 Variants for RQ2
To answer RQ2, we constructed five variants of TELPA for the ablation experiment:
TELPA that removes the backward method-invocation analysis from TELPA;
TELPA that removes the forward method-invocation analysis from TELPA;
TELPA that removes the counter-example guidance from TELPA;
TELPA that randomly selects the same number of counter-examples instead of coverage-based sampling in TELPA;
TELPA that removes the feedback process from TELPA. Specifically, the newly generated tests by the LLM, along with their coverage information, will not be added to the pool of existing tests by TELPA.
Similar to RQ1, we constructed two sets of variants based on the preceding test generation tool: Pynguin-based variants (TELPA, TELPA, TELPA, TELPA, TELPA); and CODAMOSA-based variants (TELPA, TELPA, TELPA, TELPA, TELPA).
IV-A4 Variants for RQ3
TELPA is built on top of the adopted preceding test generation tools and LLM. To better understand how this design choice will impact the performance of TELPA, in this RQ, we construct another two variants. In the first variant, named TELPA, no preceding test generation tool is used. That is, TELPA is activated from the very beginning. In the second variant, we consider a different LLM. The default LLM in TELPA is Phind-CodeLlama-34B-v2, which is relatively large-scale and may require a lot of resources to run. To make TELPA more practical, we also investigated the effectiveness of TELPA when a relatively small-scale LLM is adopted. More specifically, we employed DeepSeek-Coder-6.7B-instruct [21], which is one of the most effective LLMs among the set of LLMs with comparable (small) scales on HuggingFace. Accordingly, we call the variant TELPA. Similar to previous RQs, Pynguin and CODAMOSA based variants are created and named as TELPA and TELPA.
IV-A5 Benchmark, Metrics, and Environment
Following the existing work [4], we evaluated TELPA on the widely-studied benchmark, which consists of 486 modules from 27 Python projects. The whole benchmark contains 42897 source lines of code and 4518 methods in total.
Following the existing work on test generation [1, 4], we adopted both branch coverage and line coverage as metrics to measure the effectiveness of a test generation technique. For each module and each technique in a project, we allocated 20 minutes for test generation. All the techniques kept running until the time budget ran out, even if the coverage did not grow before that.
To avoid the influence of randomness, we repeated our experiments for three times considering the evaluation cost and used the average value of the final results. The experiments were conducted on a workstation with 128-core CPU, 504G memory, 4 NVIDIA A100 GPUs, and Ubuntu 20.04 OS.
IV-B RQ1: Effectiveness of TELPA
| Project | Pynguin | Pynguin | BasicPrompt | TELPA | CODAMOSA | CODAMOSA | BasicPrompt | TELPA |
| pysnooper | 16.95% | 18.64% | 21.47% | 29.94% | 17.07% | 17.68% | 23.00% | 31.10% |
| apimd | 39.88% | 43.30% | 55.76% | 81.31% | 7.14% | 40.06% | 9.43% | 67.39% |
| blib2to3 | 23.68% | 26.24% | 24.02% | 42.59% | 19.75% | 25.72% | 20.47% | 38.59% |
| codetiming | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
| cookiecutter | 46.28% | 55.37% | 63.64% | 65.29% | 45.45% | 57.02% | 52.16% | 66.12% |
| dataclasses_json | 16.55% | 16.55% | 28.42% | 28.06% | 17.27% | 17.63% | 29.50% | 28.42% |
| docstring_parser | 43.66% | 53.52% | 54.23% | 85.21% | 84.65% | 85.82% | 90.07% | 92.20% |
| flutes | 67.46% | 75.40% | 70.63% | 84.92% | 66.94% | 75.81% | 67.63% | 81.45% |
| flutils | 46.29% | 48.00% | 68.86% | 78.57% | 66.01% | 66.29% | 68.91% | 78.37% |
| httpie | 28.03% | 29.69% | 46.74% | 45.59% | 26.42% | 27.20% | 44.64% | 44.64% |
| isort | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% | 100.00% |
| mimesis | 77.16% | 80.50% | 79.94% | 89.69% | 68.83% | 79.11% | 72.73% | 89.42% |
| py_backwards | 28.44% | 41.74% | 36.70% | 63.76% | 30.77% | 35.75% | 36.65% | 60.63% |
| pymonet | 64.44% | 65.93% | 76.30% | 85.93% | 61.07% | 64.12% | 76.34% | 88.55% |
| pypara | 43.75% | 56.82% | 46.02% | 63.64% | 18.02% | 48.84% | 20.14% | 36.63% |
| semantic_release | 38.64% | 38.64% | 45.45% | 60.23% | 36.78% | 36.78% | 49.43% | 58.62% |
| string_utils | 83.93% | 85.71% | 93.45% | 98.21% | 64.52% | 80.00% | 72.28% | 87.88% |
| pytutils | 36.11% | 36.67% | 41.67% | 57.78% | 40.68% | 44.63% | 45.76% | 57.06% |
| sanic | 43.96% | 45.33% | 56.59% | 62.64% | 46.17% | 48.36% | 54.10% | 60.11% |
| sty | 89.13% | 91.30% | 93.48% | 95.65% | 91.30% | 94.48% | 91.30% | 95.65% |
| thefuck | 19.64% | 20.36% | 31.79% | 50.71% | 22.01% | 23.66% | 33.16% | 53.76% |
| thonny | 17.95% | 36.32% | 32.91% | 37.61% | 26.75% | 36.75% | 33.16% | 42.31% |
| tornado | 46.33% | 47.97% | 58.10% | 65.19% | 42.95% | 44.60% | 53.24% | 65.44% |
| tqdm | 14.62% | 19.23% | 19.23% | 44.62% | 35.29% | 41.18% | 37.18% | 50.00% |
| typesystem | 32.69% | 37.59% | 57.21% | 51.11% | 20.81% | 55.84% | 24.12% | 72.59% |
| youtube_dl | 15.34% | 17.03% | 15.34% | 24.90% | 20.89% | 23.54% | 20.89% | 29.11% |
| ansible | 27.87% | 29.73% | 28.19% | 38.03% | 29.76% | 31.36% | 29.84% | 37.84% |
| Average Branch Cov. | 44.77% | 48.80% | 53.56% | 64.12% | 44.72% | 51.93% | 50.23% | 63.48% |
| Average Line Cov. | 63.23% | 66.02% | 68.85% | 76.36% | 59.13% | 67.83% | 62.55% | 75.51% |
Table I presents the comparison results for TELPA vs. Pynguin vs. BasicPrompt and TELPA vs. CODAMOSA vs. BasicPrompt on each project in terms of branch coverage. To gain a better understanding of the performance of the preceding test generation step, we also show the branch coverage achieved by Pynguin and CODAMOSA right before switching to LLMs in “Pynguin” and “CODAMOSA” columns, respectively. “bn” represents “bottleneck”. The last two rows present the average branch coverage and line coverage across all projects. Due to the space limit, we do not show the line coverage for each project.
To better understand the effectiveness of TELPA and TELPA, we also measured the time spent on running Pynguin and CODAMOSA before switching to LLMs. On average across all projects, Pynguin and CODAMOSA ran for 2.57 and 3.51 minutes, respectively, before switching. That is, TELPA is activated at an early stage of the entire testing process (given a time budget of 20 minutes) regardless of taking Pynguin or CODAMOSA as the preceding test generation tool. When comparing Pynguin, Pynguin, CODAMOSA, and CODAMOSA, we found that on average, Pynguin achieves a branch coverage of 44.77% within only 2.57 minutes, but just achieves 48.80% coverage after continuing running for 17.43 minutes. Similarly, CODAMOSA achieves 44.72% branch coverage within only 3.51 minutes, but in the end just achieves 51.93% coverage after continuing running for 16.49 minutes. That is, Pynguin and CODAMOSA can quickly cover easily reachable branches, and then encounter the coverage bottleneck to some degree. This highlights the challenge posed by hard-to-cover branches and underscores the motivation behind our work.
After switching to TELPA, the branch coverage can be significantly improved according to results of TELPA and TELPA. Specifically, after reaching the testing time budget, TELPA achieves an average improvement of 31.39% over Pynguin across all projects in terms of branch coverage, and TELPA achieves an average improvement of 22.22% over CODAMOSA. In terms of line coverage, TELPA achieves an average improvement of 15.67% over Pynguin and TELPA achieves an average improvement of 11.31% over CODAMOSA. Furthermore, we conducted a paired Wilcoxon signed-rank test [22] at a significance level of 0.05 between TELPA/TELPA and Pynguin/CODAMOSA, confirming TELPA’s statistical superiority over the two state-of-the-art techniques with all p-values below 0.05.
Regardless of the preceding test generation tools (Pynguin or CODAMOSA), TELPA can achieve similar effectiveness, i.e., TELPA achieves similar branch coverage and line coverage to TELPA on average across all projects. The results demonstrate the stable effectiveness of TELPA. We conducted a paired Wilcoxon signed-rank test [22] at a significance level of 0.05 between TELPA and TELPA in terms of both branch coverage and line coverage. Both p-values exceed 0.05, demonstrating no statistically significant difference between TELPA and TELPA in improving coverage.
Compared to Pynguin, all the LLM-based techniques (including TELPA, TELPA, BasicPrompt, BasicPrompt, CODAMOSA) achieve higher branch coverage and line coverage on average across all projects. This demonstrates the effectiveness of LLMs in improving test coverage. However, different prompting methods can largely affect the effectiveness of LLMs. By comparing TELPA/TELPA with BasicPrompt/BasicPrompt, our designed prompting method specific to the challenge posed by hard-to-cover branches performs much better than the general prompting method employed for LLM-based test generation (e.g., CODAMOSA). Specifically, the average improvement of TELPA over BasicPrompt is 19.71% in terms of branch coverage and 10.92% in terms of line coverage. Similarly, the average improvement of TELPA over BasicPrompt is 26.38% in terms of branch coverage and 20.71% in terms of line coverage. This implies the importance of designing task-specific prompting, i.e., extracting useful information (rather than leveraging as much available information as possible) for improving the coverage of hard-to-cover branches in TELPA. We conducted a paired Wilcoxon signed-rank test [22] at a significance level of 0.05 and found that TELPA/TELPA significantly outperforms BasicPrompt/BasicPrompt by obtaining all p-values smaller than 0.05.
IV-C RQ2: Contribution of Each Component in TELPA
| TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | TELPA | |
| Branch Cov. | 56.32% | 58.82% | 55.44% | 58.72% | 58.60% | 64.12% | 50.79% | 53.66% | 50.31% | 51.15% | 51.58% | 63.48% |
| Line Cov. | 70.61% | 72.20% | 70.34% | 71.60% | 71.51% | 76.36% | 62.69% | 64.46% | 62.83% | 63.05% | 63.32% | 75.51% |
To investigate the contribution of each component in TELPA, we compared TELPA with a set of its variants (introduced in Section IV-A3). Table II presents the comparison results in terms of average branch coverage and average line coverage across all projects. As can be seen from Table II, both TELPA and TELPA perform better than their corresponding variants. We further performed a paired Wilcoxon signed-rank test at the significance level of 0.05 to investigate whether TELPA/TELPA significantly outperforms each variant across all projects in terms of branch coverage and line coverage. All p-values are smaller than 0.05, demonstrating the statistically significant contribution of each component in TELPA, regardless of the preceding test generation tools.
In particular, even though we removed one component from TELPA, all these variants still outperform BasicPrompt. This demonstrates the effectiveness of using part of information specific to the challenge posed by hard-to-cover branches, compared to directly integrating as much available information as possible for prompting, which also reflects the negative influence of introducing irrelevant information on LLMs.
We also investigated whether our coverage-based counter-example sampling approach contributes to the overall effectiveness of TELPA by comparing with TELPA that randomly selects counter-examples. Table II shows that TELPA outperforms TELPA regardless of the preceding test generation tools. This result further confirms the importance of using counter-examples even if they are just selected randomly. Moreover, TELPA significantly outperforms TELPA regardless of the preceding test generation tools. For example, TELPA improves TELPA by 9.42% in terms of average branch coverage achieved, while TELPA improves TELPA by 23.07%. The results demonstrate the importance of a diverse set of counter-examples, which can be well achieved by our coverage-based sampling strategy.
IV-D RQ3: TELPA’s Effectiveness with Different Configurations
| Pynguin | CODAMOSA | TELPA | TELPA | TELPA | TELPA | TELPA | |
| Branch | 48.80% | 51.93% | 53.45% | 60.27% | 64.12% | 59.67% | 63.48% |
| Line | 66.02% | 67.83% | 68.38% | 74.16% | 76.36% | 73.34% | 75.51% |
We first investigated the effectiveness of TELPA that removes the preceding test generation tool and activates TELPA from scratch. By comparing TELPA with CODAMOSA and Pynguin in Table III, we can see that the TELPA can improve CODAMOSA and Pynguin by 2.92% and 9.53% in terms of average branch coverage and 0.80% and 3.58% in terms of average line coverage. The improvements are not huge, especially compared to the differences between TELPA and CODAMOSA or between TELPA and Pynguin. This is expected, as the time needed for LLMs to generate each test cases is normally much more than that required by SBST techniques. TELPA leverages LLMs for even easy-to-cover branches, preventing it from unleashing its full potential to reach hard-to-cover branches. Instead, CODAMOSA alternately invokes Pynguin and LLMs for test generation, reducing the time needed by LLMs to a large extent, while Pynguin is also efficient as it does not rely on LLMs. Hence, within the same limited testing time, TELPA does not achieve significant coverage improvement over CODAMOSA and Pynguin. This justifies our design choice of adopting preceding test generation tools to balance effectiveness and efficiency.
By examining the data, we found that 603 methods were not covered by CODAMOSA and Pynguin at all within the testing time budget. For these methods, TELPA instead can effectively improve the coverage for 259 of them. On average across these methods, TELPA achieves 21.86% branch coverage and 39.99% line coverage, while TELPA achieves 17.14% branch coverage and 43.80% line coverage. The results further confirm the ability of TELPA in improving the coverage of hard-to-coverage branches.
We then investigated the effectiveness of TELPA and TELPA that employ DeepSeek-Coder-6.7B-instruct [21] instead of Phind-CodeLlama-34B-v2 [9]. From Table III, we found although TELPA/TELPA exhibits small performance reduction compared to TELPA/TELPA, the former still largely outperforms Pynguin/CODAMOSA. For example, TELPA/TELPA improves Pynguin/CODAMOSA by 23.50% and 14.90% in terms of average branch coverage. This demonstrates the generalizability of TELPA under different scales of LLMs to some degree.
IV-E Threats to Validity
The threats mainly lie in parameter settings in TELPA and metrics used. By default, we set the timeframe to two minutes for switching from preceding test generation tools to our LLM-based test generation, the LLM temperature to 1, and the testing time budget to 20 minutes on each software module. Both the timeframe and the testing time budget are longer than the settings in the existing studies [1, 4], demonstrating higher sufficiency of reaching coverage bottlenecks and evaluating test effectiveness. Higher temperature allows TELPA to receive responses from LLMs with higher confidence, which is commonly employed in various tasks [23, 24]. In the future, we will investigate the influence of different settings on the effectiveness of TELPA. Regarding metrics, we have employed the widely-used branch coverage and line coverage, which are also aligned with our goal (i.e., reaching hard-to-cover branches). In the future, we will use more metrics to evaluate generated tests more sufficiently, including naturalness [25] and mutation score [26].
V Discussion
Hallucination. Although TELPA can significantly improve test coverage compared to the state-of-the-art SBST and LLM-based techniques, it is still unable to reach all the hard-to-cover branches within the given testing time budget. Through our analysis, the major reason lies in the hallucination problem, namely LLMs confidently produce incorrect responses. For example, some generated tests by TELPA use non-existent parameters for invoking some methods. Currently, several approaches have been proposed to relieve the hallucination problem in various software engineering tasks, such as fine-tuning LLMs with PPO [27] or DPO [28] algorithms based on high-quality data. In the future, we can incorporate them to further improve the effectiveness of TELPA.
Orthogonality with Prompting Enhancement Methods. TELPA is a novel prompting technique for addressing the challenge of hard-to-cover branches in software testing, which improves the prompt contents specific to this task. Besides improving prompt contents, there are also some other prompting enhancement methods, including various CoT techniques (e.g., SCoT [29] and CCoT [30]) and post-processing techniques (e.g., self-repairing [20]). CoT techniques mainly elicit LLMs to produce intermediate reasoning steps based on the provided prompt for improving LLMs’ effectiveness, while post-processing techniques mainly fix invalid code produced by LLMs based on the given prompt through static analysis or providing error messages to LLMs for self-repairing. They are actually orthogonal to TELPA (improving prompt contents), thus combining them may further improve test coverage.
Generalizability. From our study, the generalizability of TELPA has been demonstrated to a certain extent by evaluating the performance on 27 projects and considering two types of preceding test generation tools (Pynguin and CODAMOSA). Besides, the idea of TELPA is not specific to Python projects (our evaluation benchmark). The involved program analyses in TELPA can be extended to the projects developed in other programming languages, as long as similar data and control flow analyses of this study can be performed on those languages. In fact, TELPA is not restricted to using a certain LLM, and many code generation-related LLMs [31, 32, 33] support program understanding across various programming languages. In the future, we will evaluate TELPA with diverse LLMs on the projects in other programming languages.
VI Related Work
We present traditional and deep-learning-based test generation techniques as related work.
The most widely-studied traditional test generation technique category is SBST, including Pynguin [1] investigated in our work. Pynguin is a state-of-the-art technique for Python projects, and there are also SBST techniques for Java projects, such as Randoop [2] and EvoSuite [3]. Randoop performs random generation, while EvoSuite leverages more advanced search algorithms to guide test generation. They have also been improved by incorporating string literals from human-written tests [34] and object construction graphs [35].
Some other techniques, instead, employ symbolic execution to improve test coverage [36, 37, 38, 39, 40]. For example, Galeotti et al. [37] integrated dynamic symbolic execution (DSE) into the Genetic Algorithm (GA) adopted by EvoSuite. Their approach mutates tests, where primitive values can influence the fitness, to cover specific corner cases without sacrificing general coverage. Braione et al. [38] and Li et al. [39] transformed path conditions into optimization problems and solved them with SBST and machine learning techniques, respectively. Baluda et al. [40] combined symbolic execution and symbolic reachability analysis to improve the effectiveness by testing rare execution conditions and eliminating infeasible branches. However, it is well-known that symbolic execution has difficulties in dealing with complex data types and objects, thus unable to address our specific challenges.
Our approach TELPA utilizes program-analysis-enhanced prompting to exploit the code comprehension ability of LLMs and facilitate the effective test generation for hard-to-cover branches. The above-mentioned techniques, while not effective for this task, can still be incorporated into TELPA as preceding testing tools and benefit from our new prompting method.
There are also several deep-learning-based test generation techniques [41, 42]. Before prompting LLMs for test generation, ATHENATEST [41] built a transformer model based on a large dataset of target methods and tests for test generation. A3Test [42] first built a pre-trained language model for assertions in a self-supervised manner based on PLBART [43], and then fine-tuned it with test generation data. Recently, some work leveraged LLMs for test generation [4, 20, 5], including CODAMOSA used in our study. For example, ChatUnitTest [20] and ChatTester [5] leveraged LLMs to generate tests with similar prompt contents used in CODAMOSA, but designed strategies (such as rule-based and LLM-based strategies) to fix invalid tests generated by LLMs.
TELPA is also a LLM-based test generation technique, which improves prompt contents with the aid of program analysis (i.e., contextual information purification specific to the challenges of hard-to-cover branches). On one hand, those test-repair strategies can be combined with TELPA to further improve its performance. On the other hand, these deep-learning-based techniques can be also used as preceding test generation tools in TELPA for orthogonal integration.
VII Conclusion
Existing test generation techniques face difficulties in covering branches involving specific complex objects and intricate inter-procedural dependencies. To address this issue, we propose a novel technique, named TELPA. TELPA leverages program analysis to assist LLMs in constructing complex objects and understanding code semantics, thus improving the test generation performance. Additionally, TELPA employs counter-example sampling and coverage-based feedback to guide LLMs to effectively and efficiently generate tests. Our experiments on 27 open-source Python projects demonstrate that TELPA significantly outperforms both the state-of-the-art SBST and LLM-based techniques.